When a hyrocarbon was completely burnt in air ,9.42g of carbon(iv)oxide and 3.87g of water were formed .find the empirical formula of hyrocarbon
1
Expert's answer
2020-04-09T13:11:20-0400

CxHy+O2=CO2+H2O


"\frac{x}{12}=\frac{9.42}{44}=2.56909091"


"\frac{y}{2}=\frac{3.87}{18}=0.43"


"=\frac{2.56909091}{12}:\frac{0.43}{1}=0.214090909:0.43=1:2"


CH2 methylene